OpenStudy (anonymous):

Which point lies on the graph of the following piecewise function? y = x + 2 if x < –1 y = x^2 if x >= –1 (0, 2) (1, –1) (–2, 4) (3, 9)

OpenStudy (hba):

why dont u plug in the points and see

OpenStudy (anonymous):

I got (-2,4) but idk if I'm right

OpenStudy (anonymous):

do you know how to read this piecewise function?

OpenStudy (anonymous):

\[f(x) = \left\{\begin{array}{rcc} x + 2 & \text{if} & x<-1 \\ x^2& \text{if} & x\geq -1 \end{array} \right. \] mean IF the input is less than \(-1\) you use the first rule \(x+2\) and add 2 to the input. but IF the input is greater than or equal to \(-1\) you use the second rule \(x^2\) and square it

OpenStudy (anonymous):

I need someone to walk me step by step to the answer, that'd be nice, I'll use this question as a refrence for my other ones